OpenStudy (anonymous):

What value of d makes the statement true? d + 12.6 = 9 A. –21.6 B. –3.6 C. 3.6 D. 21.6

OpenStudy (anonymous):

B. you must subtract 12.6 form both sides.

OpenStudy (anonymous):

basically, your solving for d d + 12.6 = 9 -- subtract 12.6 from both sides d = 9 - 12.6 d = (you can do the math)

OpenStudy (anonymous):

from* once you do that only step d is = -3.6

OpenStudy (like1234):

Solution for d+12.6=9 equation: Simplifying d + 12.6 = 9 Reorder the terms: 12.6 + d = 9 Solving 12.6 + d = 9 Solving for variable 'd'. Move all terms containing d to the left, all other terms to the right. Add '-12.6' to each side of the equation. 12.6 + -12.6 + d = 9 + -12.6 Combine like terms: 12.6 + -12.6 = 0.0 0.0 + d = 9 + -12.6 d = 9 + -12.6 Combine like terms: 9 + -12.6 = -3.6 d = -3.6 Simplifying d = -3.6 http://www.geteasysolution.com/d+12.6=9

OpenStudy (anonymous):

sorry I gave the answer away in explaining, but I figured its one easy step, you would see it anywho

OpenStudy (like1234):

the answer is B